Be sure to answer all parts.

How would you prepare 2.10

×

102 mL of a 0.726 M NaOH solution, starting with a

5.98 M

stock solution?
Volume of stock solution needed: mL
Volume of water needed: mL
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 5.98 M

M2 = 0.726 M

V2 = 210 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.726*210)/5.98

V1 = 25.5 mL

Answer: 25.5 mL

2)

volume of water = V2 - V1

= 210 mL - 25.5 mL

= 184.5 mL

Answer: 184 mL